# Sparse Training for Federated Learning with Regularized Error Correction

### v1

@@ Abstract
Sparse training schemes gain increasing attention in order to scale down the dimensionality of each client (i.e., node) transmission. Specifically, sparsification with error correction methods is a promising technique, where only important updates are sent to the parameter server (PS) and the rest are accumulated locally. While error correction methods have shown to achieve a significant sparsification level of the client-to-PS message without harming convergence, pushing sparsity further remains unresolved due to the staleness effect. In this paper, we propose a novel algorithm, dubbed Federated Learning with Accumulated Regularized Embeddings (FLARE), to overcome this challenge. FLARE presents a novel sparse training approach via accumulated pulling of the updated models with regularization on the embeddings in the FL process, providing a powerful solution to the staleness effect, and pushing sparsity to an exceptional level. The performance of FLARE is validated through extensive experiments on diverse and complex models, achieving a remarkable sparsity level (10 times and more beyond the current state-of-the-art) along with significantly improved accuracy. Additionally, an open-source software package has been developed for the benefit of researchers and developers in related fields.
@@ I-A Main Results
The Error Correction for model deltas is implemented as follows: During each iterative round of the FL process, after the client completes the local model update procedure, the client selectively identifies and transmits only those model parameter deltas deemed as the Top-$K$ in magnitude, determined by their absolute values. These selective updates are then transmitted to the PS for aggregation. Additionally, clients accumulate residuals locally, acting as corrections to counteract the Top-$K$ sparsification transmission-residuals that were not considered in the Top-$K$ are accumulated locally. Eventually, these residuals become large enough to be transmitted, ensuring that all changes are eventually sent.
@@
Second, to evaluate FLARE’s performance, we conducted extensive simulation experiments involving various ML tasks of different sizes. The experiments utilized five distinct models for different FL settings, namely Fully Connected (FC), Convolutional Neural Network (CNN), VGG 11, VGG 16, and VGG 19 models, applied to MNIST and CIFAR10 datasets. The results affirm FLARE’s superior performance compared to existing methods, demonstrating higher accuracy and sparsity levels. Specifically, FLARE achieves a sparsity level of $99.999\%$, surpassing the $99.9\%$ sparsity level of existing methods and the $99.99\%$ achieved by the current state-of-the-art. This represents a magnitude 10 times and more beyond the current state-of-the-art, accompanied by significantly improved accuracy. This remarkable advancement enables FL in bandwidth-limited networks. The robustness and efficacy of FLARE are underscored by these results, marking a significant progress in the field. The simulation results are detailed in Section IV.
@@ II System Model and Problem Statement
Notation
Description
$N$
Number of clients
$\chi^{i}$
Local dataset of client $i$
$\chi$
Global dataset
$F_{i}$
Local empirical loss for client $i$
$F$
Global empirical loss
$f$
Objective loss function
$\bar{w}^{i}_{k}$
Client’s $i$ model vector at round $k$
$\bar{w}_{k}$
Global model vector at the PS at round $k$
$\bar{A}_{k}^{i}$
Accumulator vector for client $i$ at round $k$
$E$
Number of forward-backward optimization steps
$B$
Batch size
$R$
Sparsity level
$\eta$
Learning rate
@@ III-B Algorithm Description
$\begin{array}[]{l}\displaystyle\tilde{f^{i}_{k}}(x,\bar{w};\bar{A}^{i}_{k},\bar{w_{k}})\\ \hskip 14.22636pt\displaystyle=f(x,\bar{w})\displaystyle+\tau_{k}\sum_{{\begin{subarray}{c}w\in\bar{w}\\ a\in\bar{A}^{i}_{k}\\ w_{k}\in\bar{w}_{k}\end{subarray}}}m(a)\cdot|w-(w_{k}+a)|.\end{array}$ (3)
Here,
$m(a)=\begin{cases}1,&\text{if $a>0$},\\ 0,&\text{otherwise},\end{cases}$ (4)
where $w,a,w_{k}$ are entries corresponding to $\bar{w}$, $\bar{A}^{i}$, $\bar{w}_{k}$, respectively, and the summation is over all model weights.
Algorithm 1 describes the FLARE procedure with SGD. At the beginning of each round, clients determine $\tilde{f^{i}_{k}}$ according to their accumulators. Then, each client executes a forward-backward update according to $\tilde{f^{i}_{k}}$ for $p$ steps. Then, they revert to the original loss function $f$ for the remaining steps. We next demonstrate the consecutive rounds by the algorithm step by step.
@@
                   $\bar{w}\leftarrow\bar{w}-\eta\nabla\tilde{F}^{i}_{k}(\bar{w};b)$
@@
To facilitate a clear and concise explanation of FLARE, let us consider a case of two clients, convex objective loss surface, and a two-dimensional model parameterized by two weights. For purposes of exposition, each client performs a single optimization step (i.e., $E=1$, $B$ is extremely large), and uses a sparsity level of $50\%$ (i.e., $R=50\%$) in the distributed SGD procedure. This means that at every transmission from clients to the PS, the dimension with large update is sent to the PS, where the small one is kept at the clients. The loss surfaces for each client, $F_{1}(w;\chi_{1}),F_{2}(w;\chi_{2})$, can be seen in Fig. 1 colored by red and green, respectively, and the global objective is colored by black. We refer to the two-dimensional model weights that constitute $\bar{w}=(w_{1},w_{2})$ as horizontal and vertical. Also, recall that $w,a,w_{i}$ are the entries of $\bar{w},\bar{A},\bar{w}_{i}$, respectively, and when written at the same equation, referring to the same weight location of the model, without using additional indexing.
The description is provided below in four steps, accompanied by Figs. 0(a)-0(d).
@@
    $\begin{array}[]{l}\displaystyle\tilde{f^{i}_{0}}(x,\bar{w};\bar{A}^{i}_{0},\bar{w_{0}})\\ \hskip 14.22636pt\displaystyle=f(x,\bar{w})\displaystyle+\tau\sum_{{\begin{subarray}{c}w\in\bar{w}\\ a\in\bar{A}^{i}_{0}\\ w_{0}\in\bar{w}_{0}\end{subarray}}}m(a)\cdot|w-(w_{k}+a)|.\end{array}$ (5)
@@
$\tau||\bar{w}-(\bar{w}_{1}+\bar{A}^{i}_{1})||_{1}=\tau\sum_{{\begin{subarray}{c}w\in\bar{w}\\ a\in\bar{A}^{i}_{1}\\ w_{1}\in\bar{w}_{1}\end{subarray}}}|w-(w_{1}+a)|.$ (7)
Minimizing (7) allows for the compensation of staled updates at each optimization step utilizing residual information, without any communication or computational costs. While the minimization of (7) influences both weights, we intended to pull only staled updates, i.e., the vertical update in our example at this stage. Pulling the horizontal weights is unwanted and may impede learning. As all of the horizontal accumulated content was just released, its accumulated value equals zero, leaving the horizontal weight to be regularized according to $|w-w_{1}|$ which serves no purpose but to slow convergence [^1] The accumulator can be served as an indicator of how staled is an update at a specific round. The larger its accumulator content, the more staled it is.. To tackle this issue, a masking operation is performed by detecting those up-to-date weights according to their accumulated values by the function $m(a)$, defined by (4). The expression $a>0$ indicates the weight is staled and should be pulled. With this, (8) below pulls only staled updates. In addition, a new hyperparameter $\tau$ is set to control the pulling and the loss objective ratio:
$\tau\sum_{{\begin{subarray}{c}w\in\bar{w}\\ a\in\bar{A}^{i}_{1}\\ w_{1}\in\bar{w}_{1}\end{subarray}}}m(a)\cdot|w-(w_{0}+a)|.$ (8)
@@
    $\begin{array}[]{l}\displaystyle\tilde{f^{i}_{1}}(x,\bar{w};\bar{A}^{i}_{1},\bar{w_{1}})\\ \hskip 14.22636pt\displaystyle=f(x,\bar{w})\displaystyle+\tau\sum_{{\begin{subarray}{c}w\in\bar{w}\\ a\in\bar{A}^{i}_{1}\\ w_{1}\in\bar{w}_{1}\end{subarray}}}m(a)\cdot|w-(w_{1}+a)|,\end{array}$ (9)
@@ III-C Further Insights and Observations of FLARE Algorithm
Advantages of accumulated pulling and masking in FLARE: The accumulated pulling implemented by FLARE helps each client to compensate for an unsent update at each round, and not only at the round the update is released, giving a direct solution to the staleness effect. Secondly, it creates a chain reaction: More accurate accumulation in the second round aids each client in generating a new model with greater precision in the third round, enhancing the third accumulation, which, in turn, improves the fourth, and so forth. Lastly, masking can be done in various ways. To mitigate the regularization effect, masking can be defined more generally using a threshold $t$:
$m(a)=\begin{cases}1,&\text{if $a>t$},\\ 0,&\text{otherwise}.\end{cases}$ (12)
The larger $t$, the less weights are affected by FLARE. The threshold $t$ can be set for example as the average of all accumulator values.
@@
To address this issue, we utilize the FLARE regularization term only for the first $p$ optimization steps in each round. In each round, only the initial $p$ steps are optimized according to $\tilde{f^{i}_{k}}(x,\bar{w};\bar{A}^{i}_{k},\bar{w}^{i}_{k})$, while the subsequent steps are optimized according to $f(x,\bar{w})$. This approach allows only the initial steps to correct for stale updates while preventing the rest from being influenced by any pulling that could lead to contamination. Additionally, we introduce a decay in the parameter $\tau$ as a function of $k$ with an exponential decay rule, setting $\tau_{k+1}=\tau_{k}/{c}$. These adjustments enhance the applicability of FLARE for multiple steps, mitigating the impact of contaminated accumulators on the regularization term. These observations are demonstrated in the experiments outlined in Section IV.
IV Experiments

### v2

@@ Abstract
Sparse training schemes gain increasing attention in order to scale down the dimensionality of each client (i.e., node) transmission. Specifically, sparsification with error correction methods is a promising technique, where only important updates are sent to the parameter server (PS) and the rest are accumulated locally. While error correction methods have shown to achieve a significant sparsification level of the client-to-PS message without harming convergence, pushing sparsity further remains unresolved due to the staleness effect. In this paper, we propose a novel algorithm, dubbed Federated Learning with Accumulated Regularized Embeddings (FLARE), to overcome this challenge. FLARE presents a novel sparse training approach via accumulated pulling of the updated models with regularization on the embeddings in the FL process, providing a powerful solution to the staleness effect, and pushing sparsity to an exceptional level. Our theoretical analysis demonstrates that FLARE’s regularized error feedback achieves significant improvements in scalability with sparsity parameter. The empirical performance of FLARE is validated through extensive experiments on diverse and complex models, achieving a remarkable sparsity level (10 times and more beyond the current state-of-the-art) along with significantly improved accuracy. Additionally, an open-source software package has been developed for the benefit of researchers and developers in related fields.
@@ I-A Main Results
The Error Correction for model deltas is implemented as follows: During each iterative round of the FL process, after the client completes the local model update procedure, the client selectively identifies and transmits only those model parameter deltas deemed as the Top-$K$ in magnitude, determined by their absolute values. These selective updates are then transmitted to the PS for aggregation. Additionally, clients accumulate residuals locally, acting as corrections to counteract the Top-$K$ sparsification. That is, residuals initially omitted from the Top-$K$ selection are accumulated locally. Over time, these residuals become large enough to be transmitted, ensuring that all changes are eventually sent.
@@
Second, for performance analysis, we conducted both theoretical and empirical evaluations to assess the efficiency of FLARE. The theoretical convergence analysis demonstrates that FLARE’s regularized error feedback not only matches state-of-the-art techniques in terms of convergence rate with time, but also achieves significant improvements in scalability with respect to the sparsity parameter. In terms of empirical evaluations, we conducted extensive simulation experiments involving various ML tasks of different sizes. The experiments utilized six distinct models for different FL settings, namely Fully Connected (FC), Convolutional Neural Network (CNN), VGG 11, VGG 16, VGG 19 and GRU models, applied to MNIST, CIFAR10, and ”The Complete Works of William Shakespeare” datasets. The results affirm FLARE’s superior performance compared to existing methods, demonstrating higher accuracy and sparsity levels. Specifically, FLARE achieves a sparsity level of $99.999\%$, surpassing the $99.9\%$ sparsity level of existing methods and the $99.99\%$ achieved by the current state-of-the-art. This represents a magnitude 10 times and more beyond the current state-of-the-art, accompanied by significantly improved accuracy. This remarkable advancement enables FL in bandwidth-limited networks. The robustness and efficacy of FLARE are underscored by these results, marking a significant progress in the field. The simulation results are detailed in Section V.
@@ II System Model and Problem Statement
Notation Description
$N$ Number of clients
$d$ Weight dimension
$T$ Number of rounds
$k$ Round index
$\chi^{i}$ Local dataset of client $i$
$\chi$ Global dataset
$F_{i}$ Local empirical loss for client $i$
$F$ Global empirical loss
$f$ Objective loss function
$\bar{w}^{i}_{k}$ Client’s $i$ model vector at round $k$
$\bar{w}_{k}$ Global model vector at the PS at round $k$
$\bar{A}_{k}^{i}$ Accumulator vector for client $i$ at round $k$
$E$ Number of optimization steps
$B$ Batch size
$R,\delta$ Sparsity percent, sparsity ratio
$\gamma$ Learning rate
$\tau,\tau_{k}$ FLARE regularization coefficient
$a_{0}$ FLARE masking coefficient
$c$ decay constant of $\tau$
$p$ Number of steps for FLARE objective
@@ III-B Algorithm Description
$\begin{array}[]{l}\displaystyle\tilde{f^{i}_{k}}(x,\bar{w};\bar{A}^{i}_{k},\bar{w_{k}})\vspace{0.2cm}\\ \hskip 14.22636pt\displaystyle=f(x,\bar{w})\displaystyle+\tau_{k}\sum^{d}_{j=1}m(a_{j})\cdot|w_{j}-(w_{k,j}+a_{j})|.\end{array}$ (3)
Here, $w_{j},a_{j},w_{k,j}$ are $j$-th entries of $\bar{w}$, $\bar{A}_{k}^{i}$, $\bar{w}_{k}$, respectively (where we omit complex indexing notations when the context is clear), summed over all model weights, $\tau_{k}$ is a regularization constant possibly varies with time, and
$m(a)=\begin{cases}1,&\text{if $|a|>a_{0}$},\\ 0,&\text{otherwise}.\end{cases}$ (4)
Algorithm 1 describes the FLARE procedure with SGD. At the beginning of each round, clients determine $\tilde{f^{i}_{k}}$ according to their accumulators. Then, each client executes a forward-backward update according to $\tilde{f^{i}_{k}}$ for $p$ steps. Then, they revert to the original loss function $f$ for the remaining steps. Note that regularization in this description is based on the $L_{1}$ norm, but it can also be applied using the $L_{2}$ norm. Our theoretical results are developed specifically with respect to the $L_{2}$ norm. We next demonstrate the consecutive rounds by the algorithm step by step.
@@
                   $\bar{w}\leftarrow\bar{w}-\gamma\nabla\tilde{F}^{i}_{k}(\bar{w};b)$
@@
To facilitate a clear and concise intuitive explanation of FLARE, let us consider a case of two clients, convex objective loss surface, and a two-dimensional model parameterized by two weights. For purposes of exposition, each client performs a single optimization step (i.e., $E=1$, $B=\infty$), and uses a sparsity level of $50\%$ (i.e., $R=50\%$) in the distributed SGD procedure with $a_{0}$ set to zero. This means that at every transmission from clients to the PS, the dimension with large update is sent to the PS, where the small one is kept at the clients. The loss surfaces for each client, $F_{1}(w;\chi_{1}),F_{2}(w;\chi_{2})$, can be seen in Fig. 1 colored by red and green, respectively, and the global objective is colored by black. We refer to the two-dimensional model weights that constitute $\bar{w}=(w_{1},w_{2})$ as horizontal and vertical.
The description is provided below in four steps, accompanied by Figs. 1(a)-1(d).
@@
    $\begin{array}[]{l}\displaystyle\tilde{f^{i}_{0}}(x,\bar{w};\bar{A}^{i}_{0},\bar{w_{0}})\vspace{0.2cm}\\ \hskip 14.22636pt\displaystyle=f(x,\bar{w})\displaystyle+\tau_{0}\sum^{2}_{j=1}m(a_{j})\cdot|w_{j}-(w_{k,j}+a_{j})|.\end{array}$ (5)
@@
$\tau_{1}||\bar{w}-(\bar{w}_{1}+\bar{A}^{i}_{1})||^{2}_{1}=\tau_{1}\sum^{2}_{j=1}|w_{j}-(w_{1,j}+a_{j})|$ (7)
Minimizing (7) allows for the compensation of staled updates at each optimization step utilizing residual information, without any communication or computational costs. While the minimization of (7) influences both weights, we intended to pull only staled updates, i.e., the vertical update in our example at this stage. Pulling the horizontal weights is unwanted and may impede learning. As all of the horizontal accumulated content was just released, its accumulated value equals zero, leaving the horizontal weight to be regularized according to $|w-w_{1}|$ which serves no purpose but to slow convergence [^1] The accumulator can be served as an indicator of how staled is an update at a specific round. The larger its accumulator content, the more staled it is.. To tackle this issue, a masking operation is performed by detecting those up-to-date weights according to their accumulated values by the function $m(a)$, defined by (4). The expression $|a|>0$ indicates the weight is staled and should be pulled. With this, pulling is done only for staled updates. In addition, a new hyper parameter $\tau_{k}$ is set to control the pulling and the loss objective ratio.
@@
    $\begin{array}[]{l}\displaystyle\tilde{f^{i}_{1}}(x,\bar{w};\bar{A}^{i}_{1},\bar{w_{1}})\vspace{0.2cm}\\ \hskip 14.22636pt\displaystyle=f(x,\bar{w})\displaystyle+\tau_{1}\sum^{2}_{j=1}m(a_{j})\cdot|w_{j}-(w_{1,j}+a_{j})|,\end{array}$ (8)
@@ III-C Further Insights and Observations of FLARE Algorithm
Advantages of accumulated pulling and masking in FLARE: The accumulated pulling implemented by FLARE helps each client to compensate for an unsent update at each round, and not only at the round the update is released, giving a direct solution to the staleness effect. Secondly, it creates a chain reaction: More accurate accumulation in the second round aids each client in generating a new model with greater precision in the third round, enhancing the third accumulation, which, in turn, improves the fourth, and so forth. Lastly, masking can be done in various ways. To mitigate the regularization effect, masking can be defined more generally using a threshold $a_{0}$ as in (4). The larger $a_{0}$, the less weights are affected by FLARE. The threshold $a_{0}$ can be set for example as the average of all accumulator values.
@@
To address this issue, we utilize the FLARE regularization term only for the first $p$ optimization steps in each round. In each round, only the initial $p$ steps are optimized according to $\tilde{f^{i}_{k}}(x,\bar{w};\bar{A}^{i}_{k},\bar{w}^{i}_{k})$, while the subsequent steps are optimized according to $f(x,\bar{w})$. This approach allows only the initial steps to correct for stale updates while preventing the rest from being influenced by any pulling that could lead to contamination. Additionally, we introduce a decay in the parameter $\tau$ as a function of $k$ with an exponential decay rule, setting $\tau_{k+1}=\tau_{k}/{c}$. These adjustments enhance the applicability of FLARE for multiple steps, mitigating the impact of contaminated accumulators on the regularization term. These observations are demonstrated in the experiments outlined in Section V.
IV Convergence Analysis
In this section we analyze the convergence of FLARE. Let $\delta$ be the sparsification ratio, $10<\delta<$ 1, and $C(\bar{w})$ be the Top-$K$ sparsification operator. Consistent with the assumptions used in [37, 38] and related studies, we introduce the following assumptions for our analysis: Assumption 1 pertains to the compressor, Assumption 2 addresses smoothness, and Assumption 3 concerns the bounded stochastic gradient moment.
Assumption 1.
A Compression operator $C:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a $\delta$ approximation for $10<\delta<$ 1 if for all $\bar{w}\in\mathbb{R}^{d}$, the following inequality holds: $||C(\bar{w})-\bar{w}||^{2}\leq(1-\delta)||\bar{w}||^{2}.$
Assumption 2.
Function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $L$-smooth if there exists a constant $L>0$ such that for all $\bar{x},\bar{y}\in\mathbb{R}^{d}$, the following inequality holds: $f(\bar{y})\leq f(\bar{x})+\nabla f(\bar{x})^{T}(\bar{y}-\bar{x})+\frac{L}{2}\|\bar{y}-\bar{x}\|^{2}.$
Assumption 3.
For any $\bar{w}\in\mathbb{R}^{d}$, the stochastic gradient fulfills: $\mathbb{E}||\nabla f(\bar{w})||^{2}\leq\sigma^{2}.$
We begin by reviewing the analytic convergence performance achieved by the EC and sparsification method in [37]. As detailed in [38, 37], the error correction and sparsification methods are analytically validated for a single client. These methods, however, provide significant benefits when extended to multi-client settings, greatly enhancing their overall performance. It is also developed for a single optimization step at each iteration (i.e., FedSGD setting). In our paper, we adopt similar assumptions for the analysis of our proposed method. Specifically, we set $N=1,E=1,B=\infty$, $a_{0}=0$ in FLARE algorithm. Throughout the analysis, we refer to the loss function by $F$, and the regularized loss function under FLARE by $\tilde{F}$, where $\tilde{F}(\bar{w};\bar{w}_{t},\bar{A}_{t})=F(\bar{w})+\frac{\tau}{2}||\bar{w}-(\bar{w}_{t}+\bar{A}_{t})||^{2}.$
Consider the case where $F$ is convex, and Assumptions 1 and 3 hold. Let $\bar{w}_{0}$ be the initial model, $\{\bar{w}_{t}\}_{t>0}$ be the sequence of updated model iterations produced by EC method [37], with step size $\gamma$, and let $\bar{w}^{avg}_{T}=\frac{1}{T}\sum^{T}_{t=0}{\bar{w}_{t}}$ be the averaged update. Then, the error under EC is bounded by [37]:
$\mathbb{E}[F(\bar{w}^{avg}_{T})]-F^{*}\leq\dfrac{||\bar{w}_{0}-w^{*}||^{2}}{2\gamma(T+1)}+\gamma\sigma^{2}\left(\dfrac{1}{2}+\dfrac{2\sqrt{1-\delta}}{\delta}\right),$ (11)
where $w^{*}$ is the minimizer of $F$, and $F^{*}=F(w^{*})$.
Secondly, in the non-convex scenario, assuming Assumptions 1, 2, and 3 hold, it can be shown that [37]:
$\min_{t\in[T]}\mathbb{E}||\nabla F(\bar{w}_{t})||^{2}\leq$ $\dfrac{2(F(\bar{w}_{0})-F(w^{*}))}{\gamma(T+1)}+L\gamma\sigma^{2}+$
$\hskip 43.05542pt\dfrac{4\gamma^{2}L^{2}(1-\delta)\sigma^{2}}{\delta^{2}}.$ (12)
Setting $\gamma$ on the order of $\frac{1}{\sqrt{T+1}}$ achieves the best convergence rate order $O(1/\sqrt{T})$ with $T$ as will be demonstrated later.
Setting $\delta=1$ results in the convergence rate of SGD[38, 37]. Furthermore, by choosing $\gamma$ on the order of $\frac{1}{\sqrt{T+1}}$ while keeping $\delta$ fixed, the method asymptotically approaches the performance of SGD as $T$ increases. However, in practical scenarios for finite $T$, the influence of $\delta$-related terms becomes significant. In particular, reducing $\delta$ to increase the sparsification level can degrade the performance of EC through these terms. Specifically, when examining the scaling rate with respect to the sparsification ratio $\delta$ while keeping $T$ finite, it becomes evident that the error bound in (11) scales with a rate of $1/\delta$, and the gradient moment in (12) scales with a rate of $1/\delta^{2}$, as $\delta$ approaches zero. This degradation is notably pronounced under extreme sparsification settings, where the performance of EC significantly deteriorates.
In the following theorems, we demonstrate that FLARE offers substantial performance enhancements, particularly in mitigating the performance degradation associated with increased sparsification levels. Specifically, we show below that for the error bound in the convex case scales with a rate of $1/\sqrt{\delta}$, and the gradient moment in the non-convex case scales with a rate of $1/\delta$, as $\delta$ approaches zero. Furthermore, FLARE achieves the same order of $1/\sqrt{T}$ with $T$ as EC.
Figure 2: An illustration of the function $p(\tau)$.
Theorem 1.
Consider the case where Assumption 1, 3 hold, and the loss function $F$ is convex. Denote the model iterates of FLARE by $\{\bar{w}_{t}\}_{t>0}$, with step size $\gamma$, and let $\bar{w}^{avg}_{T}=\frac{1}{T}\sum^{T}_{t=0}{\bar{w}_{t}}$. Denote the minimizer of $\tilde{F}$ by $w^{*}$,and let $\tilde{F}^{*}=\tilde{F}(w^{*})$. Then,
$\mathbb{E}[\tilde{F}(\bar{w}^{avg}_{T})]-\tilde{F}^{*}\leq\dfrac{||\bar{w}_{0}-w^{*}||^{2}}{2\gamma(T+1)}+$
$\hskip 86.11084pt\gamma\sigma^{2}\left(\dfrac{1}{2}+\dfrac{2\sqrt{1-\delta}}{\delta\sqrt{1+p(\tau)}}\right),$ (13)
where
$p(\tau)=\dfrac{2\gamma}{\delta}(2-\delta)\tau+\dfrac{2\gamma^{2}}{\delta}(\dfrac{\delta}{2}-1)\tau^{2}.$ (14)
The proof can be found in the Appendix.
Next, we set $\gamma=\dfrac{a}{\sqrt{T+1}}$ as done in [37], with $a=\nicefrac{{||\bar{w}_{0}-w^{*}||}}{{\sigma}}$, to achieve the best convergence rate with $T$. Then, the error bound under FLARE (13) becomes:
$\mathbb{E}[\tilde{F}(\bar{w}^{avg}_{T})]-\tilde{F}^{*}\leq$ $\frac{\left\|\bar{w}_{0}-w^{*}\right\|}{2\sqrt{T+1}}\sigma\left(2+\frac{4\sqrt{1-\delta}}{\sqrt{\delta^{2}(1+p(\tau))}}\right).$ (15)
For comparison, the error bound under EC (11) becomes:
$\mathbb{E}[F(\bar{w}^{avg}_{T})]-F^{*}\leq\dfrac{||\bar{w}_{0}-w^{*}||}{2\sqrt{T+1}}\sigma\left(2+\dfrac{4\sqrt{1-\delta}}{\sqrt{\delta^{2}}}\right).$ (16)
As illustrated in Fig. 2, the function $p(\tau)$ is parabolic with $p(\tau)>0$ for $0<\tau<\frac{2}{\gamma}$ with maximum at $\nicefrac{{(2-\delta)}}{{\delta}}$. For $\tau=0$, $p(0)=0$, which aligns with the established convergence rate of the EC method as shown in (11). However, our approach improves upon the EC method by leveraging the FLARE regularization coefficient to reduce the error scaling. Specifically, the positive factor $\sqrt{1+p(\tau)}$ helps to mitigate the scaling with $\delta$. By choosing $\tau$ such that $p(\tau)$ scales as $1/\delta$, FLARE achieves an error bound that scales as $1/\sqrt{\delta}$, in contrast to the $1/\delta$ scaling under EC. Note that as $\delta$ approaches zero, for a finite $T$, $\tau$ can be arbitrarily small, by setting: $\tau=1/(\gamma\cdot C)$ for any fixed $C>1$, yielding $p(\tau)=\frac{2C-1}{C^{2}}\cdot\frac{2-\delta}{\delta}$ scales with $1/\delta$. Additionally, in terms of the scaling order with $T$, FLARE maintains the same order of $1/\sqrt{T}$ as achieved by EC.
Theorem 2.
Consider the case where Assumption 1,2 and 3 hold. Denote the model iterates of FLARE by $\{\bar{w}_{t}\}_{t>0}$, with step size $\gamma$. Denote the minimizer of $\tilde{F}$ by $w^{*}$. Then,
$\min_{t\in[T]}\mathbb{E}||\nabla\tilde{F}(\bar{w}_{t+1})||^{2}\leq\dfrac{2(F(\bar{w}_{0})-F(w^{*}))}{\gamma(T+1)}+L\gamma\sigma^{2}+$
$\hskip 21.52771pt\dfrac{2}{\gamma}\left(\dfrac{L^{2}\gamma}{2}+r(\tau)\right)\dfrac{4\gamma^{2}(1-\delta)\sigma^{2}}{\delta^{2}}\cdot\dfrac{1}{1+p(\tau)},$ (17)
where
$r(\tau)=\left(\dfrac{\gamma}{2}+\dfrac{L\gamma^{2}}{2}\right)\tau^{2}.$ (18)
The proof can be found in the Appendix.
Note that $r(\tau)>0$ for $\tau>0$. Again, we are aligned with the known rate of EC (12) for $\tau=0$.
Next, we set $\gamma=\dfrac{1}{\sqrt{T+1}}$ as done in [37] to achieve the best convergence rate with $T$. Then, the gradient moment bound under FLARE (17) becomes:
$\min_{t\in[T]}\mathbb{E}||\nabla\tilde{F}(\bar{w}_{t+1})||^{2}\leq\dfrac{2(F(w_{0})-F(w^{*}))+L\sigma^{2}}{\sqrt{T+1}}+$
$\dfrac{4L^{2}(1-\delta)\sigma^{2}}{\delta^{2}(T+1)}\cdot\dfrac{1}{1+p(\tau)}+\dfrac{4(1-\delta)\sigma^{2}}{(\nicefrac{{\delta}}{{\tau}})^{2}(T+1)}\cdot\dfrac{1}{1+p(\tau)}$
$\hskip 42.67912pt+\dfrac{2}{(\nicefrac{{\delta}}{{\tau}})^{2}}\dfrac{4(1-\delta)\sigma^{2}}{\sqrt{T+1}(T+1)}\cdot\dfrac{1}{1+p(\tau)}.$ (19)
For comparison, the gradient moment bound under EC (12) becomes:
$\min_{t\in[T]}\mathbb{E}||\nabla F($ $\bar{w}_{t+1})||^{2}\leq\dfrac{2(F(\bar{w}_{0})-F(w^{*}))+L\sigma^{2}}{\sqrt{T+1}}$
$\hskip 56.9055pt+\dfrac{4L^{2}(1-\delta)\sigma^{2}}{\delta^{2}(T+1)}.$ (20)
For $\tau=0$, $p(0)=0$, and the results are aligned with the known rate of EC (11) in this case as well. However, it can be seen again that our approach improves upon the EC method by leveraging the FLARE regularization coefficient to reduce the error scaling. By choosing $\tau$ such that $p(\tau)$ scales as $1/\delta$, FLARE achieves an error bound that scales as $1/\delta$, in contrast to the $1/\delta^{2}$ scaling under EC. Additionally, in terms of the scaling order with $T$, FLARE maintains the same order of $1/\sqrt{T}$ as achieved by EC.
V Experiments